13. Solve the following systems of differential equations. r = x +y, y = 3.x – y X = 8x + 10y, y = -5x – 7y
Answer #1

det (A-17=0 29-0 71=2,-21 Xtv = Get () +68443). ego. W X - AY, A 18 10 dit (A-17) = 0 3 1 2 1 - 6 - 2 A = 18 det (A-1 T-07 1²